OpenStudy (anonymous):

Differentiate y=(x^2+2)(cube root (x^2+2))

OpenStudy (e.mccormick):

Looks like product rule and a chain.

OpenStudy (anonymous):

i just learn chain rule today so i am a little unsure of where to use it

OpenStudy (e.mccormick):

AH. OK. Fist, let me write this out in LaTeX.

OpenStudy (e.mccormick):

\(y=(x^2+2)(\sqrt[3]{x^2+2})\) OK. for the product rule part you have (u)(v). That is not too bad. But when you do the product rule you end up with needing the primes of each. That is where the chain comes in.

OpenStudy (e.mccormick):

The cahin rule states that whan you take a derivative of something that contains a derivative, you do a little multiplication: \[\frac{dy}{dx}f(g(x))=f'(g(x))g'(x)\]What that means is take the derivative of the outer function, which is the cube root in the right most part of yours, and leave the inner function alone. Then, multiply that times the derivitive of the iner part.

OpenStudy (anonymous):

alright let me try to solve it

OpenStudy (e.mccormick):

So when you need v' you need \(\frac{dy}{xd}\sqrt[3]{x^2+2}\) Start by rewriting that as a fractional exponent. \[(\sqrt[3]{x^2+2})'\Rightarrow (x^2+2)^{1/3} \] Then it should be easer to do that part.

OpenStudy (anonymous):

okay so thats where i use the chain rule right

OpenStudy (e.mccormick):

Yah.

OpenStudy (anonymous):

so for that i got 1/3(x^2+2)^-2/3(2x)

OpenStudy (e.mccormick):

And that is for the v' part. So your uv=u'v+uv' it is just that last v' part.

OpenStudy (raden):

hint :